OpenStudy (this):

helppp plzzzz -5k^2+10k-13=-8 a,0;two imaginary solution b,0;0ne real solution c,c,200;two real solution d,0;two real solution

OpenStudy (anonymous):

start with \(5k^2-10k+5=0\) then maybe divide by 5 and get \[k^2-2k+1=0\]

OpenStudy (anonymous):

this is a perfect square, namely \((k-1)^2=0\) and so there is only one real solution

OpenStudy (this):

thank you @satellite73

OpenStudy (anonymous):

yw